Calculate the pH of the resulting solution when the original solution is mixed with 35.0 mL of 0.975 M NaOH solution


1
Expert's answer
2021-04-29T07:17:49-0400

An erlenmeyer flask contains 50.0 mL of an aqueous solution of 0.425 M acetic acid (CH3COOH). Given that acetic acid has a Ka = 1.75×10-5. Calculate the pH of the resulting solution when the original solution is mixed with 35.0 mL of 0.975 M NaOH solution.


Solution:

Vsolution = 50.0 mL + 35.0 mL = 85.0 mL

C(CH3COOH) = 0.425 M × (50.0 mL / 85.0 mL) = 0.25 M

C(NaOH) = 0.975 M × (35.0 mL / 85.0 mL) = 0.40 M


The balanced chemical equation:

CH3COOH + NaOH → CH3COONa + H2O

ICE Table:



NaOH >> CH3COONa

So, pH is determined by the concentration of excess NaOH

NaOH → Na+ + OH-

[OH-] = C(NaOH) = 0.15 M (according to the ICE table)

pOH = -log[OH-] = -log(0.15) = 0.82

pH = 14 - pOH = 14 - 0.82 = 13.18

pH = 13.18


Answer: pH = 13.18